OpenStudy (anonymous):

MEDAL!!! (a). What is the period? (b). What is the amplitude? (c). What is the frequency? (d). How is this graph related to its parent graph y=sinx and y=cosx

OpenStudy (anonymous):

y=2sin(2xpi)+1

OpenStudy (anonymous):

a.) period for a sin graph is = 2pi/(b) so here it would be 2pi/(2)= pi since the 2s cancel so your period is pi b.)amplitude is the number in front of your trigonometric function, in this case its 2 c.) not entirely sure what you mean by this d.) vertical stretch of 2, left pi, up 1

OpenStudy (anonymous):

Erm, I don't know this stuff

OpenStudy (anonymous):

@shubham.bagrecha

OpenStudy (anonymous):

is that 2 * pi in y ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

no is it 2pix

OpenStudy (anonymous):

you know period of sin ?

OpenStudy (anonymous):

2pi

OpenStudy (anonymous):

yes If a periodic function has period 'T' and there is a function like F=f(kx), then period of F will be T/k.

OpenStudy (anonymous):

So, here , T=2pi k(coefficient of x) = 2pi therefore period of given function is T/k = 1.

OpenStudy (anonymous):

understood ?

OpenStudy (anonymous):

so the period of y=2sin(2Pix)+1 is 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok I get that so far

OpenStudy (anonymous):

I know the amplitude is 2

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

frequency= 1?

OpenStudy (anonymous):

is it a physics problem as i don't know frequency in maths ?

OpenStudy (anonymous):

if yes frequency=1 is correct

OpenStudy (anonymous):

no this is precalc/trig.

OpenStudy (anonymous):

but frequency is just T/1

OpenStudy (anonymous):

are you 100% sure the period is 1? Just making sure lol

OpenStudy (anonymous):

frequency=1/T=1

OpenStudy (anonymous):

Sorry that is correct. Is the period 1 for sure though because that affects the frequency answer

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

frequency is 1/t not t/1 so you were right. The period is 1 though, right? You are 100% sure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so I get all that. What about D)

OpenStudy (anonymous):

d) i can only say that all these have same curves (shapes)

OpenStudy (anonymous):

D) wants to know about the phase shifts and stuff. I can talk about the different periods and amplitudes